OpenStudy (anonymous):

Please help me solve this. -2/5+p=-4/15

OpenStudy (anonymous):

In fraction form

OpenStudy (anonymous):

-2/5 + p = -4/15 Change -2/5 into -6/15 by multiplying by 3/3. -6/15 + p = -4/15 (now subtract -6/15 from both sides.) p = -4/15 - (-6/15) (double negative make a positive) p = -4/15 + 6/15 p = 2/15

OpenStudy (anonymous):

Does that make sense?